\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{100}

\bibitem{Cimatti}
A.~{Cimatti}, F.~{Fraternali} and C.~{Nipoti}, \emph{{Introduction to Galaxy
  Formation and Evolution: From Primordial Gas to Present-Day Galaxies}}
  (2019).

\bibitem{Hubble1929}
E.~{Hubble}, \emph{{A Relation between Distance and Radial Velocity among
  Extra-Galactic Nebulae}},
  \href{https://doi.org/10.1073/pnas.15.3.168}{\emph{Proceedings of the
  National Academy of Science} {\bfseries 15} (1929) 168}.

\bibitem{Riess2020}
A.G.~{Riess}, \emph{{The expansion of the Universe is faster than expected}},
  \href{https://doi.org/10.1038/s42254-019-0137-0}{\emph{Nature Reviews
  Physics} {\bfseries 2} (2020) 10}
  [\href{https://arxiv.org/abs/2001.03624}{{\ttfamily 2001.03624}}].

\bibitem{Bullock2017}
J.S.~{Bullock} and M.~{Boylan-Kolchin}, \emph{{Small-Scale Challenges to the
  {\ensuremath{\Lambda}}CDM Paradigm}},
  \href{https://doi.org/10.1146/annurev-astro-091916-055313}{\emph{\araa}
  {\bfseries 55} (2017) 343}
  [\href{https://arxiv.org/abs/1707.04256}{{\ttfamily 1707.04256}}].

\bibitem{Padovani2017}
P.~{Padovani}, D.M.~{Alexander}, R.J.~{Assef}, B.~{De Marco}, P.~{Giommi},
  R.C.~{Hickox} et~al., \emph{{Active galactic nuclei: what's in a name?}},
  \href{https://doi.org/10.1007/s00159-017-0102-9}{\emph{\aapr} {\bfseries 25}
  (2017) 2} [\href{https://arxiv.org/abs/1707.07134}{{\ttfamily 1707.07134}}].

\bibitem{Treu2010}
T.~{Treu}, \emph{{Strong Lensing by Galaxies}},
  \href{https://doi.org/10.1146/annurev-astro-081309-130924}{\emph{\araa}
  {\bfseries 48} (2010) 87} [\href{https://arxiv.org/abs/1003.5567}{{\ttfamily
  1003.5567}}].

\bibitem{Banados2018}
E.~{Ba{\~n}ados}, B.P.~{Venemans}, C.~{Mazzucchelli}, E.P.~{Farina},
  F.~{Walter}, F.~{Wang} et~al., \emph{{An 800-million-solar-mass black hole in
  a significantly neutral Universe at a redshift of 7.5}},
  \href{https://doi.org/10.1038/nature25180}{\emph{\nat} {\bfseries 553} (2018)
  473} [\href{https://arxiv.org/abs/1712.01860}{{\ttfamily 1712.01860}}].

\bibitem{Belladitta2022}
S.~{Belladitta}, A.~{Caccianiga}, A.~{Diana}, A.~{Moretti}, P.~{Severgnini},
  M.~{Pedani} et~al., \emph{{Central engine of the highest redshift blazar}},
  \href{https://doi.org/10.1051/0004-6361/202142335}{\emph{\aap} {\bfseries
  660} (2022) A74} [\href{https://arxiv.org/abs/2201.08863}{{\ttfamily
  2201.08863}}].

\bibitem{Piana2021}
O.~{Piana}, P.~{Dayal}, M.~{Volonteri} and T.R.~{Choudhury}, \emph{{The mass
  assembly of high-redshift black holes}},
  \href{https://doi.org/10.1093/mnras/staa3363}{\emph{\mnras} {\bfseries 500}
  (2021) 2146} [\href{https://arxiv.org/abs/2009.13505}{{\ttfamily
  2009.13505}}].

\bibitem{Chiaberge2015}
M.~{Chiaberge}, R.~{Gilli}, J.M.~{Lotz} and C.~{Norman}, \emph{{Radio Loud AGNs
  are Mergers}}, \href{https://doi.org/10.1088/0004-637X/806/2/147}{\emph{\apj}
  {\bfseries 806} (2015) 147}
  [\href{https://arxiv.org/abs/1505.07419}{{\ttfamily 1505.07419}}].

\bibitem{Gao2020}
F.~{Gao}, L.~{Wang}, W.J.~{Pearson}, Y.A.~{Gordon}, B.W.~{Holwerda},
  A.M.~{Hopkins} et~al., \emph{{Mergers trigger active galactic nuclei out to z
  {\ensuremath{\sim}} 0.6}},
  \href{https://doi.org/10.1051/0004-6361/201937178}{\emph{\aap} {\bfseries
  637} (2020) A94} [\href{https://arxiv.org/abs/2004.00680}{{\ttfamily
  2004.00680}}].

\bibitem{Whitler2023}
L.~{Whitler}, R.~{Endsley}, D.P.~{Stark}, M.~{Topping}, Z.~{Chen} and
  S.~{Charlot}, \emph{{On the ages of bright galaxies 500 Myr after the big
  bang: insights into star formation activity at z {\ensuremath{\gtrsim}} 15
  with JWST}}, \href{https://doi.org/10.1093/mnras/stac3535}{\emph{\mnras}
  {\bfseries 519} (2023) 157}
  [\href{https://arxiv.org/abs/2208.01599}{{\ttfamily 2208.01599}}].

\bibitem{Pacucci2022}
F.~{Pacucci}, P.~{Dayal}, Y.~{Harikane}, A.K.~{Inoue} and A.~{Loeb}, \emph{{Are
  the newly-discovered z 13 drop-out sources starburst galaxies or quasars?}},
  \href{https://doi.org/10.1093/mnrasl/slac035}{\emph{\mnras} {\bfseries 514}
  (2022) L6} [\href{https://arxiv.org/abs/2201.00823}{{\ttfamily 2201.00823}}].

\bibitem{Volonteri2023}
M.~{Volonteri}, M.~{Habouzit} and M.~{Colpi}, \emph{{What if young z > 9 JWST
  galaxies hosted massive black holes?}},
  \href{https://doi.org/10.1093/mnras/stad499}{\emph{\mnras} {\bfseries 521}
  (2023) 241} [\href{https://arxiv.org/abs/2212.04710}{{\ttfamily
  2212.04710}}].

\bibitem{Momjian2021}
E.~{Momjian}, E.~{Ba{\~n}ados}, C.L.~{Carilli}, F.~{Walter} and
  C.~{Mazzucchelli}, \emph{{Resolving the Radio Emission from the Quasar
  P172+18 at z = 6.82}},
  \href{https://doi.org/10.3847/1538-3881/abe6ae}{\emph{\aj} {\bfseries 161}
  (2021) 207} [\href{https://arxiv.org/abs/2103.03481}{{\ttfamily
  2103.03481}}].

\bibitem{Momjian2008}
E.~{Momjian}, C.L.~{Carilli} and I.D.~{McGreer}, \emph{{Very Large Array and
  Very Long Baseline Array Observations of the Highest Redshift Radio-Loud QSO
  J1427+3312 at Z = 6.12}},
  \href{https://doi.org/10.1088/0004-6256/136/1/344}{\emph{\aj} {\bfseries 136}
  (2008) 344} [\href{https://arxiv.org/abs/0805.2897}{{\ttfamily 0805.2897}}].

\bibitem{Frey2010}
S.~{Frey}, Z.~{Paragi}, L.I.~{Gurvits}, D.~{Cseh} and K.{\'E}.~{Gab{\'a}nyi},
  \emph{{High-resolution images of five radio quasars at early cosmological
  epochs}}, \href{https://doi.org/10.1051/0004-6361/201015554}{\emph{\aap}
  {\bfseries 524} (2010) A83}
  [\href{https://arxiv.org/abs/1009.5023}{{\ttfamily 1009.5023}}].

\bibitem{Frey2011}
S.~{Frey}, Z.~{Paragi}, L.I.~{Gurvits}, K.{\'E}.~{Gab{\'a}nyi} and D.~{Cseh},
  \emph{{Into the central 10 pc of the most distant known radio quasar. VLBI
  imaging observations of J1429+5447 at z = 6.21}},
  \href{https://doi.org/10.1051/0004-6361/201117341}{\emph{\aap} {\bfseries
  531} (2011) L5} [\href{https://arxiv.org/abs/1106.0717}{{\ttfamily
  1106.0717}}].

\bibitem{Frey2013}
S.~{Frey}, J.O.~{Fogasy}, Z.~{Paragi} and L.I.~{Gurvits}, \emph{{On the Doppler
  boosting in the compact radio jet of the distant blazar J1026+2542 at z =
  5.3}}, \href{https://doi.org/10.1093/mnras/stt249}{\emph{\mnras} {\bfseries
  431} (2013) 1314} [\href{https://arxiv.org/abs/1302.2209}{{\ttfamily
  1302.2209}}].

\bibitem{Cao2014}
H.M.~{Cao}, S.~{Frey}, L.I.~{Gurvits}, J.~{Yang}, X.Y.~{Hong}, Z.~{Paragi}
  et~al., \emph{{VLBI observations of the radio quasar J2228+0110 at z = 5.95
  and other field sources in multiple-phase-centre mode}},
  \href{https://doi.org/10.1051/0004-6361/201323328}{\emph{\aap} {\bfseries
  563} (2014) A111} [\href{https://arxiv.org/abs/1402.2353}{{\ttfamily
  1402.2353}}].

\bibitem{Gabanyi2015}
K.E.~{Gabanyi}, D.~{Cseh}, S.~{Frey}, Z.~{Paragi}, L.I.~{Gurvits}, T.~{An}
  et~al., \emph{{VLBI observation of the newly discovered z = 5.18 quasar SDSS
  J0131-0321.}}, \href{https://doi.org/10.1093/mnrasl/slv046}{\emph{\mnras}
  {\bfseries 450} (2015) L57}
  [\href{https://arxiv.org/abs/1503.06593}{{\ttfamily 1503.06593}}].

\bibitem{Zhang2017}
Y.~{Zhang}, T.~{An}, S.~{Frey}, K.{\'E}.~{Gab{\'a}nyi}, Z.~{Paragi},
  L.I.~{Gurvits} et~al., \emph{{J0906+6930: a radio-loud quasar in the early
  Universe}}, \href{https://doi.org/10.1093/mnras/stx392}{\emph{\mnras}
  {\bfseries 468} (2017) 69}
  [\href{https://arxiv.org/abs/1702.03925}{{\ttfamily 1702.03925}}].

\bibitem{Cao2017}
H.M.~{Cao}, S.~{Frey}, K.{\'E}.~{Gab{\'a}nyi}, Z.~{Paragi}, J.~{Yang},
  D.~{Cseh} et~al., \emph{{VLBI observations of four radio quasars at z > 4:
  blazars or not?}}, \href{https://doi.org/10.1093/mnras/stx160}{\emph{\mnras}
  {\bfseries 467} (2017) 950}
  [\href{https://arxiv.org/abs/1701.04760}{{\ttfamily 1701.04760}}].

\bibitem{Wang2017}
R.~{Wang}, E.~{Momjian}, C.L.~{Carilli}, X.-B.~{Wu}, X.~{Fan}, F.~{Walter}
  et~al., \emph{{Milliarcsecond Imaging of the Radio Emission from the Quasar
  with the Most Massive Black Hole at Reionization}},
  \href{https://doi.org/10.3847/2041-8213/835/2/L20}{\emph{\apjl} {\bfseries
  835} (2017) L20} [\href{https://arxiv.org/abs/1701.03260}{{\ttfamily
  1701.03260}}].

\bibitem{Momjian2018}
E.~{Momjian}, C.L.~{Carilli}, E.~{Ba{\~n}ados}, F.~{Walter} and
  B.P.~{Venemans}, \emph{{Resolving the Powerful Radio-loud Quasar at z
  {\ensuremath{\sim}} 6}},
  \href{https://doi.org/10.3847/1538-4357/aac76f}{\emph{\apj} {\bfseries 861}
  (2018) 86} [\href{https://arxiv.org/abs/1807.02560}{{\ttfamily 1807.02560}}].

\bibitem{Frey2018}
S.~{Frey}, O.~{Titov}, A.E.~{Melnikov}, P.~{de Vicente} and F.~{Shu},
  \emph{{High-resolution radio imaging of two luminous quasars beyond redshift
  4.5}}, \href{https://doi.org/10.1051/0004-6361/201832771}{\emph{\aap}
  {\bfseries 618} (2018) A68}
  [\href{https://arxiv.org/abs/1807.06837}{{\ttfamily 1807.06837}}].

\bibitem{Perger2019}
K.~{Perger}, S.~{Frey}, K.{\'E}.~{Gab{\'a}nyi} and L.V.~{T{\'o}th},
  \emph{{Unveiling the weak radio quasar population at z{\ensuremath{\geq}}4}},
  \href{https://doi.org/10.1093/mnras/stz2723}{\emph{\mnras} {\bfseries 490}
  (2019) 2542} [\href{https://arxiv.org/abs/1909.11012}{{\ttfamily
  1909.11012}}].

\bibitem{Spingola2020}
C.~{Spingola}, D.~{Dallacasa}, S.~{Belladitta}, A.~{Caccianiga},
  M.~{Giroletti}, A.~{Moretti} et~al., \emph{{Parsec-scale properties of the
  radio brightest jetted AGN at z > 6}},
  \href{https://doi.org/10.1051/0004-6361/202039458}{\emph{\aap} {\bfseries
  643} (2020) L12} [\href{https://arxiv.org/abs/2010.11193}{{\ttfamily
  2010.11193}}].

\bibitem{An2020}
T.~{An}, P.~{Mohan}, Y.~{Zhang}, S.~{Frey}, J.~{Yang}, K.{\'E}.~{Gab{\'a}nyi}
  et~al., \emph{{Evolving parsec-scale radio structure in the most distant
  blazar known}},
  \href{https://doi.org/10.1038/s41467-019-14093-2}{\emph{Nature
  Communications} {\bfseries 11} (2020) 143}
  [\href{https://arxiv.org/abs/2001.02809}{{\ttfamily 2001.02809}}].

\bibitem{Zhang2021}
Y.~{Zhang}, T.~{An}, S.~{Frey}, X.~{Yang}, M.~{Krezinger}, O.~{Titov} et~al.,
  \emph{{J2102+6015: a young radio source at z = 4.575}},
  \href{https://doi.org/10.1093/mnras/stab2289}{\emph{\mnras} {\bfseries 507}
  (2021) 3736} [\href{https://arxiv.org/abs/2108.02142}{{\ttfamily
  2108.02142}}].

\bibitem{Gabanyi2021}
K.{\'E}.~{Gab{\'a}nyi}, S.~{Frey}, T.~{An}, H.~{Cao}, Z.~{Paragi},
  L.I.~{Gurvits} et~al., \emph{{A small radio galaxy at z = 4.026}},
  \href{https://doi.org/10.1002/asna.20210057}{\emph{Astronomische Nachrichten}
  {\bfseries 342} (2021) 1092}
  [\href{https://arxiv.org/abs/2110.10964}{{\ttfamily 2110.10964}}].

\bibitem{Perger2021}
K.~{Perger}, S.~{Frey}, D.A.~{Schwartz}, K.{\'E}.~{Gab{\'a}nyi}, L.I.~{Gurvits}
  and Z.~{Paragi}, \emph{{Multi-scale Radio and X-Ray Structure of the
  High-redshift Quasar PMN J0909+0354}},
  \href{https://doi.org/10.3847/1538-4357/ac0144}{\emph{\apj} {\bfseries 915}
  (2021) 98} [\href{https://arxiv.org/abs/2105.06307}{{\ttfamily 2105.06307}}].

\bibitem{Zhang2022}
Y.~{Zhang}, T.~{An}, A.~{Wang}, S.~{Frey}, L.I.~{Gurvits},
  K.{\'E}.~{Gab{\'a}nyi} et~al., \emph{{VLBI observations of VIK
  J2318{\ensuremath{-}}3113, a quasar at z = 6.44}},
  \href{https://doi.org/10.1051/0004-6361/202243785}{\emph{\aap} {\bfseries
  662} (2022) L2} [\href{https://arxiv.org/abs/2205.05859}{{\ttfamily
  2205.05859}}].

\bibitem{Krezinger2022}
M.~{Krezinger}, K.~{Perger}, K.{\'E}.~{Gab{\'a}nyi}, S.~{Frey}, L.I.~{Gurvits},
  Z.~{Paragi} et~al., \emph{{Radio-loud Quasars above Redshift 4: Very Long
  Baseline Interferometry (VLBI) Imaging of an Extended Sample}},
  \href{https://doi.org/10.3847/1538-4365/ac63b8}{\emph{\apjs} {\bfseries 260}
  (2022) 49} [\href{https://arxiv.org/abs/2204.02114}{{\ttfamily 2204.02114}}].

\bibitem{Walter2014}
F.~{Walter}, R.~{Decarli}, M.~{Sargent}, C.~{Carilli}, M.~{Dickinson},
  D.~{Riechers} et~al., \emph{{A Molecular Line Scan in the Hubble Deep Field
  North: Constraints on the CO Luminosity Function and the Cosmic H$_{2}$
  Density}}, \href{https://doi.org/10.1088/0004-637X/782/2/79}{\emph{\apj}
  {\bfseries 782} (2014) 79} [\href{https://arxiv.org/abs/1312.6365}{{\ttfamily
  1312.6365}}].

\bibitem{Aravena2016a}
M.~{Aravena}, J.S.~{Spilker}, M.~{Bethermin}, M.~{Bothwell}, S.C.~{Chapman},
  C.~{de Breuck} et~al., \emph{{A survey of the cold molecular gas in
  gravitationally lensed star-forming galaxies at z > 2}},
  \href{https://doi.org/10.1093/mnras/stw275}{\emph{\mnras} {\bfseries 457}
  (2016) 4406} [\href{https://arxiv.org/abs/1602.00652}{{\ttfamily
  1602.00652}}].

\bibitem{Aravena2016b}
M.~{Aravena}, R.~{Decarli}, F.~{Walter}, R.~{Bouwens}, P.A.~{Oesch},
  C.L.~{Carilli} et~al., \emph{{The ALMA Spectroscopic Survey in the Hubble
  Ultra Deep Field: Search for [CII] Line and Dust Emission in 6}},
  \href{https://doi.org/10.3847/1538-4357/833/1/71}{\emph{\apj} {\bfseries 833}
  (2016) 71} [\href{https://arxiv.org/abs/1607.06772}{{\ttfamily 1607.06772}}].

\bibitem{Decarli2016}
R.~{Decarli}, F.~{Walter}, M.~{Aravena}, C.~{Carilli}, R.~{Bouwens}, E.~{da
  Cunha} et~al., \emph{{ALMA Spectroscopic Survey in the Hubble Ultra Deep
  Field: CO Luminosity Functions and the Evolution of the Cosmic Density of
  Molecular Gas}},
  \href{https://doi.org/10.3847/1538-4357/833/1/69}{\emph{\apj} {\bfseries 833}
  (2016) 69} [\href{https://arxiv.org/abs/1607.06770}{{\ttfamily 1607.06770}}].

\bibitem{Riechers2019}
D.A.~{Riechers}, R.~{Pavesi}, C.E.~{Sharon}, J.A.~{Hodge}, R.~{Decarli},
  F.~{Walter} et~al., \emph{{COLDz: Shape of the CO Luminosity Function at High
  Redshift and the Cold Gas History of the Universe}},
  \href{https://doi.org/10.3847/1538-4357/aafc27}{\emph{\apj} {\bfseries 872}
  (2019) 7} [\href{https://arxiv.org/abs/1808.04371}{{\ttfamily 1808.04371}}].

\bibitem{Spingola2020_gas}
C.~{Spingola}, J.P.~{McKean}, S.~{Vegetti}, D.~{Powell}, M.W.~{Auger},
  L.V.E.~{Koopmans} et~al., \emph{{SHARP - VI. Evidence for CO (1-0) molecular
  gas extended on kpc-scales in AGN star-forming galaxies at high redshift}},
  \href{https://doi.org/10.1093/mnras/staa1342}{\emph{\mnras} {\bfseries 495}
  (2020) 2387} [\href{https://arxiv.org/abs/1905.06363}{{\ttfamily
  1905.06363}}].

\bibitem{Bischetti2021}
M.~{Bischetti}, C.~{Feruglio}, E.~{Piconcelli}, F.~{Duras},
  M.~{P{\'e}rez-Torres}, R.~{Herrero} et~al., \emph{{The WISSH quasars project.
  IX. Cold gas content and environment of luminous QSOs at z
  {\ensuremath{\sim}} 2.4-4.7}},
  \href{https://doi.org/10.1051/0004-6361/202039057}{\emph{\aap} {\bfseries
  645} (2021) A33} [\href{https://arxiv.org/abs/2009.01112}{{\ttfamily
  2009.01112}}].

\bibitem{DiMatteo2008}
T.~{Di Matteo}, J.~{Colberg}, V.~{Springel}, L.~{Hernquist} and D.~{Sijacki},
  \emph{{Direct Cosmological Simulations of the Growth of Black Holes and
  Galaxies}}, \href{https://doi.org/10.1086/524921}{\emph{\apj} {\bfseries 676}
  (2008) 33} [\href{https://arxiv.org/abs/0705.2269}{{\ttfamily 0705.2269}}].

\bibitem{Dave2020}
R.~{Dav{\'e}}, R.A.~{Crain}, A.R.H.~{Stevens}, D.~{Narayanan}, A.~{Saintonge},
  B.~{Catinella} et~al., \emph{{Galaxy cold gas contents in modern cosmological
  hydrodynamic simulations}},
  \href{https://doi.org/10.1093/mnras/staa1894}{\emph{\mnras} {\bfseries 497}
  (2020) 146} [\href{https://arxiv.org/abs/2002.07226}{{\ttfamily
  2002.07226}}].

\bibitem{Khusanova2022}
Y.~{Khusanova}, E.~{Ba{\~n}ados}, C.~{Mazzucchelli}, S.~{Rojas-Ruiz},
  E.~{Momjian}, F.~{Walter} et~al., \emph{{The [CII] and FIR properties of z >
  6 radio-loud quasars}},
  \href{https://doi.org/10.1051/0004-6361/202243660}{\emph{\aap} {\bfseries
  664} (2022) A39} [\href{https://arxiv.org/abs/2204.08973}{{\ttfamily
  2204.08973}}].

\bibitem{Riechers2008a}
D.A.~{Riechers}, F.~{Walter}, B.J.~{Brewer}, C.L.~{Carilli}, G.F.~{Lewis},
  F.~{Bertoldi} et~al., \emph{{A Molecular Einstein Ring at z = 4.12: Imaging
  the Dynamics of a Quasar Host Galaxy Through a Cosmic Lens}},
  \href{https://doi.org/10.1086/591434}{\emph{\apj} {\bfseries 686} (2008) 851}
  [\href{https://arxiv.org/abs/0806.4616}{{\ttfamily 0806.4616}}].

\bibitem{Riechers2008b}
D.A.~{Riechers}, F.~{Walter}, C.L.~{Carilli}, F.~{Bertoldi} and E.~{Momjian},
  \emph{{Formation of a Quasar Host Galaxy through a Wet Merger 1.4 Billion
  Years after the Big Bang}},
  \href{https://doi.org/10.1086/592834}{\emph{\apjl} {\bfseries 686} (2008) L9}
  [\href{https://arxiv.org/abs/0808.3774}{{\ttfamily 0808.3774}}].

\bibitem{Riechers2011}
D.A.~{Riechers}, C.L.~{Carilli}, R.J.~{Maddalena}, J.~{Hodge}, A.I.~{Harris},
  A.J.~{Baker} et~al., \emph{{CO(J = 1{\textrightarrow}0) in z > 2 Quasar Host
  Galaxies: No Evidence for Extended Molecular Gas Reservoirs}},
  \href{https://doi.org/10.1088/2041-8205/739/1/L32}{\emph{\apjl} {\bfseries
  739} (2011) L32} [\href{https://arxiv.org/abs/1106.2553}{{\ttfamily
  1106.2553}}].

\bibitem{Deane2013}
R.P.~{Deane}, I.~{Heywood}, S.~{Rawlings} and P.J.~{Marshall}, \emph{{The
  preferentially magnified active nucleus in IRAS F10214+4724 - II. Spatially
  resolved cold molecular gas}},
  \href{https://doi.org/10.1093/mnras/stt957}{\emph{\mnras} {\bfseries 434}
  (2013) 23} [\href{https://arxiv.org/abs/1307.6565}{{\ttfamily 1307.6565}}].

\bibitem{Paraficz2018}
D.~{Paraficz}, M.~{Rybak}, J.P.~{McKean}, S.~{Vegetti}, D.~{Sluse},
  F.~{Courbin} et~al., \emph{{ALMA view of RX J1131-1231: Sub-kpc CO (2-1)
  mapping of a molecular disk in a lensed star-forming quasar host galaxy}},
  \href{https://doi.org/10.1051/0004-6361/201731250}{\emph{\aap} {\bfseries
  613} (2018) A34} [\href{https://arxiv.org/abs/1705.09931}{{\ttfamily
  1705.09931}}].

\bibitem{Begelman1980}
M.C.~{Begelman}, R.D.~{Blandford} and M.J.~{Rees}, \emph{{Massive black hole
  binaries in active galactic nuclei}},
  \href{https://doi.org/10.1038/287307a0}{\emph{Nature} {\bfseries 287} (1980)
  307}.

\bibitem{DiMatteo2005}
T.~{Di Matteo}, V.~{Springel} and L.~{Hernquist}, \emph{{Energy input from
  quasars regulates the growth and activity of black holes and their host
  galaxies}}, \href{https://doi.org/10.1038/nature03335}{\emph{\nat} {\bfseries
  433} (2005) 604} [\href{https://arxiv.org/abs/astro-ph/0502199}{{\ttfamily
  astro-ph/0502199}}].

\bibitem{Volonteri2021}
M.~{Volonteri}, M.~{Habouzit} and M.~{Colpi}, \emph{{The origins of massive
  black holes}}, \href{https://doi.org/10.1038/s42254-021-00364-9}{\emph{Nature
  Reviews Physics} {\bfseries 3} (2021) 732}
  [\href{https://arxiv.org/abs/2110.10175}{{\ttfamily 2110.10175}}].

\bibitem{BurkeSpolaor2019}
S.~{Burke-Spolaor}, S.R.~{Taylor}, M.~{Charisi}, T.~{Dolch}, J.S.~{Hazboun},
  A.M.~{Holgado} et~al., \emph{{The astrophysics of nanohertz gravitational
  waves}}, \href{https://doi.org/10.1007/s00159-019-0115-7}{\emph{\aapr}
  {\bfseries 27} (2019) 5} [\href{https://arxiv.org/abs/1811.08826}{{\ttfamily
  1811.08826}}].

\bibitem{Colpi2019}
M.~{Colpi}, K.~{Holley-Bockelmann}, T.~{Bogdanovic}, P.~{Natarajan},
  J.~{Bellovary}, A.~{Sesana} et~al., \emph{{Astro2020 science white paper: The
  gravitational wave view of massive black holes}}, {\emph{arXiv e-prints}
  (2019) arXiv:1903.06867} [\href{https://arxiv.org/abs/1903.06867}{{\ttfamily
  1903.06867}}].

\bibitem{DeRosa2019}
A.~{De Rosa}, C.~{Vignali}, T.~{Bogdanovi{\'c}}, P.R.~{Capelo}, M.~{Charisi},
  M.~{Dotti} et~al., \emph{{The quest for dual and binary supermassive black
  holes: A multi-messenger view}},
  \href{https://doi.org/10.1016/j.newar.2020.101525}{\emph{New Astronomy
  Reviews} {\bfseries 86} (2019) 101525}
  [\href{https://arxiv.org/abs/2001.06293}{{\ttfamily 2001.06293}}].

\bibitem{Rodriguez2006}
C.~{Rodriguez}, G.B.~{Taylor}, R.T.~{Zavala}, A.B.~{Peck}, L.K.~{Pollack} and
  R.W.~{Romani}, \emph{{A Compact Supermassive Binary Black Hole System}},
  \href{https://doi.org/10.1086/504825}{\emph{\apj} {\bfseries 646} (2006) 49}
  [\href{https://arxiv.org/abs/astro-ph/0604042}{{\ttfamily
  astro-ph/0604042}}].

\bibitem{Bondi2010}
M.~{Bondi} and M.A.~{P{\'e}rez-Torres}, \emph{{VLBI Detection of an Active
  Galactic Nucleus Pair in the Binary Black Hole Candidate SDSS J1536+0441}},
  \href{https://doi.org/10.1088/2041-8205/714/2/L271}{\emph{\apjl} {\bfseries
  714} (2010) L271} [\href{https://arxiv.org/abs/1003.4849}{{\ttfamily
  1003.4849}}].

\bibitem{Burke-Spolaor2011}
S.~{Burke-Spolaor}, \emph{{A radio Census of binary supermassive black holes}},
  \href{https://doi.org/10.1111/j.1365-2966.2010.17586.x}{\emph{\mnras}
  {\bfseries 410} (2011) 2113}
  [\href{https://arxiv.org/abs/1008.4382}{{\ttfamily 1008.4382}}].

\bibitem{An2013}
T.~{An}, Z.~{Paragi}, S.~{Frey}, T.~{Xiao}, W.A.~{Baan}, S.~{Komossa} et~al.,
  \emph{{The radio structure of 3C 316, a galaxy with double-peaked narrow
  optical emission lines}},
  \href{https://doi.org/10.1093/mnras/stt801}{\emph{\mnras} {\bfseries 433}
  (2013) 1161} [\href{https://arxiv.org/abs/1305.1509}{{\ttfamily 1305.1509}}].

\bibitem{Deane2014}
R.P.~{Deane}, Z.~{Paragi}, M.J.~{Jarvis}, M.~{Coriat}, G.~{Bernardi},
  R.P.~{Fender} et~al., \emph{{A close-pair binary in a distant triple
  supermassive black hole system}},
  \href{https://doi.org/10.1038/nature13454}{\emph{\nat} {\bfseries 511} (2014)
  57} [\href{https://arxiv.org/abs/1406.6365}{{\ttfamily 1406.6365}}].

\bibitem{Gabanyi2016}
K.{\'E}.~{Gab{\'a}nyi}, T.~{An}, S.~{Frey}, S.~{Komossa}, Z.~{Paragi},
  X.Y.~{Hong} et~al., \emph{{Four Dual AGN Candidates Observed with the VLBA}},
  \href{https://doi.org/10.3847/0004-637X/826/2/106}{\emph{\apj} {\bfseries
  826} (2016) 106} [\href{https://arxiv.org/abs/1605.09188}{{\ttfamily
  1605.09188}}].

\bibitem{Spingola2019_mg2016}
C.~{Spingola}, J.P.~{McKean}, D.~{Massari} and L.V.E.~{Koopmans}, \emph{{Proper
  motion in lensed radio jets at redshift 3: A possible dual super-massive
  black hole system in the early Universe}},
  \href{https://doi.org/10.1051/0004-6361/201935427}{\emph{\aap} {\bfseries
  630} (2019) A108} [\href{https://arxiv.org/abs/1908.11756}{{\ttfamily
  1908.11756}}].

\bibitem{Schwartz2021}
D.~{Schwartz}, C.~{Spingola} and A.~{Barnacka}, \emph{{Resolving Complex Inner
  X-Ray Structure of the Gravitationally Lensed AGN MG B2016+112}},
  \href{https://doi.org/10.3847/1538-4357/ac0909}{\emph{\apj} {\bfseries 917}
  (2021) 26} [\href{https://arxiv.org/abs/2103.08537}{{\ttfamily 2103.08537}}].

\bibitem{Mangat2021}
C.S.~{Mangat}, J.P.~{McKean}, R.~{Brilenkov}, P.~{Hartley}, H.R.~{Stacey},
  S.~{Vegetti} et~al., \emph{{PS J1721+8842: a gravitationally lensed dual AGN
  system at redshift 2.37 with two radio components}},
  \href{https://doi.org/10.1093/mnrasl/slab106}{\emph{\mnras} {\bfseries 508}
  (2021) L64} [\href{https://arxiv.org/abs/2109.03253}{{\ttfamily
  2109.03253}}].

\bibitem{Nair2005}
S.~{Nair}, C.~{Jin} and M.A.~{Garrett}, \emph{{Helical jet in the
  gravitationally lensed blazar PKS1830-211}},
  \href{https://doi.org/10.1111/j.1365-2966.2005.09355.x}{\emph{\mnras}
  {\bfseries 362} (2005) 1157}
  [\href{https://arxiv.org/abs/astro-ph/0507015}{{\ttfamily
  astro-ph/0507015}}].

\bibitem{Spingola2022}
C.~{Spingola}, D.~{Schwartz} and A.~{Barnacka}, \emph{{Milliarcsecond X-Ray
  Astrometry to Resolve Inner Regions of AGN at z > 1 Using Gravitational
  Lensing}}, \href{https://doi.org/10.3847/1538-4357/ac68eb}{\emph{\apj}
  {\bfseries 931} (2022) 68}
  [\href{https://arxiv.org/abs/2203.04245}{{\ttfamily 2203.04245}}].

\bibitem{Delpopolo2021}
A.~{Del Popolo} and M.~{Le Delliou}, \emph{{Review of Solutions to the
  Cusp-Core Problem of the {\ensuremath{\Lambda}}CDM Model}},
  \href{https://doi.org/10.3390/galaxies9040123}{\emph{Galaxies} {\bfseries 9}
  (2021) 123}.

\bibitem{Lovell2012}
M.R.~{Lovell}, V.~{Eke}, C.S.~{Frenk}, L.~{Gao}, A.~{Jenkins}, T.~{Theuns}
  et~al., \emph{{The haloes of bright satellite galaxies in a warm dark matter
  universe}},
  \href{https://doi.org/10.1111/j.1365-2966.2011.20200.x}{\emph{\mnras}
  {\bfseries 420} (2012) 2318}
  [\href{https://arxiv.org/abs/1104.2929}{{\ttfamily 1104.2929}}].

\bibitem{Hartley2019}
P.~{Hartley}, N.~{Jackson}, D.~{Sluse}, H.R.~{Stacey} and H.~{Vives-Arias},
  \emph{{Strong lensing reveals jets in a sub-microJy radio-quiet quasar}},
  \href{https://doi.org/10.1093/mnras/stz510}{\emph{\mnras} {\bfseries 485}
  (2019) 3009} [\href{https://arxiv.org/abs/1901.05791}{{\ttfamily
  1901.05791}}].

\bibitem{Spingola2018}
C.~{Spingola}, J.P.~{McKean}, M.W.~{Auger}, C.D.~{Fassnacht},
  L.V.E.~{Koopmans}, D.J.~{Lagattuta} et~al., \emph{{SHARP - V. Modelling
  gravitationally lensed radio arcs imaged with global VLBI observations}},
  \href{https://doi.org/10.1093/mnras/sty1326}{\emph{\mnras} {\bfseries 478}
  (2018) 4816} [\href{https://arxiv.org/abs/1807.05566}{{\ttfamily
  1807.05566}}].

\bibitem{Biggs2004}
A.D.~{Biggs}, I.W.A.~{Browne}, N.J.~{Jackson}, T.~{York}, M.A.~{Norbury},
  J.P.~{McKean} et~al., \emph{{Radio, optical and infrared observations of
  CLASS B0128+437}},
  \href{https://doi.org/10.1111/j.1365-2966.2004.07701.x}{\emph{\mnras}
  {\bfseries 350} (2004) 949}
  [\href{https://arxiv.org/abs/astro-ph/0402128}{{\ttfamily
  astro-ph/0402128}}].

\bibitem{McKean2007}
J.P.~{McKean}, L.V.E.~{Koopmans}, C.E.~{Flack}, C.D.~{Fassnacht},
  D.~{Thompson}, K.~{Matthews} et~al., \emph{{High-resolution imaging of the
  anomalous flux ratio gravitational lens system CLASS B2045+265: dark or
  luminous satellites?}},
  \href{https://doi.org/10.1111/j.1365-2966.2007.11744.x}{\emph{\mnras}
  {\bfseries 378} (2007) 109}
  [\href{https://arxiv.org/abs/astro-ph/0611215}{{\ttfamily
  astro-ph/0611215}}].

\bibitem{More2009}
A.~{More}, J.P.~{McKean}, S.~{More}, R.W.~{Porcas}, L.V.E.~{Koopmans} and
  M.A.~{Garrett}, \emph{{The role of luminous substructure in the gravitational
  lens system MG 2016+112}},
  \href{https://doi.org/10.1111/j.1365-2966.2008.14342.x}{\emph{\mnras}
  {\bfseries 394} (2009) 174}
  [\href{https://arxiv.org/abs/0810.5341}{{\ttfamily 0810.5341}}].

\bibitem{Hsueh2016}
J.W.~{Hsueh}, C.D.~{Fassnacht}, S.~{Vegetti}, J.P.~{McKean}, C.~{Spingola},
  M.W.~{Auger} et~al., \emph{{SHARP - II. Mass structure in strong lenses is
  not necessarily dark matter substructure: a flux ratio anomaly from an
  edge-on disc in B1555+375}},
  \href{https://doi.org/10.1093/mnrasl/slw146}{\emph{\mnras} {\bfseries 463}
  (2016) L51} [\href{https://arxiv.org/abs/1601.01671}{{\ttfamily
  1601.01671}}].

\bibitem{Hsueh2017}
J.W.~{Hsueh}, L.~{Oldham}, C.~{Spingola}, S.~{Vegetti}, C.D.~{Fassnacht},
  M.W.~{Auger} et~al., \emph{{SHARP - IV. An apparent flux-ratio anomaly
  resolved by the edge-on disc in B0712+472}},
  \href{https://doi.org/10.1093/mnras/stx1082}{\emph{\mnras} {\bfseries 469}
  (2017) 3713} [\href{https://arxiv.org/abs/1701.06575}{{\ttfamily
  1701.06575}}].

\bibitem{Stacey2020}
H.R.~{Stacey}, A.~{Lafontaine} and J.P.~{McKean}, \emph{{Smoke on the water: CO
  and H$_{2}$O in a circumnuclear disc around a quasar at redshift 2.64}},
  \href{https://doi.org/10.1093/mnras/staa494}{\emph{\mnras} {\bfseries 493}
  (2020) 5290} [\href{https://arxiv.org/abs/2002.07186}{{\ttfamily
  2002.07186}}].

\bibitem{Powell2021}
D.~{Powell}, S.~{Vegetti}, J.P.~{McKean}, C.~{Spingola}, F.~{Rizzo} and
  H.R.~{Stacey}, \emph{{A novel approach to visibility-space modelling of
  interferometric gravitational lens observations at high angular resolution}},
  \href{https://doi.org/10.1093/mnras/staa2740}{\emph{\mnras} {\bfseries 501}
  (2021) 515} [\href{https://arxiv.org/abs/2005.03609}{{\ttfamily
  2005.03609}}].

\bibitem{Powell2022}
D.M.~{Powell}, S.~{Vegetti}, J.P.~{McKean}, C.~{Spingola}, H.R.~{Stacey} and
  C.D.~{Fassnacht}, \emph{{A lensed radio jet at milliarcsecond resolution I:
  Bayesian comparison of parametric lens models}},
  \href{https://doi.org/10.1093/mnras/stac2350}{\emph{\mnras} {\bfseries 516}
  (2022) 1808} [\href{https://arxiv.org/abs/2207.03375}{{\ttfamily
  2207.03375}}].

\bibitem{Gurvits1993}
L.I.~{Gurvits}, \emph{{Milli-arcsecond structure of AGN at different redshifts
  (Review)}},  in \emph{Sub-arcsecond Radio Astronomy}, R.J.~{Davis} and
  R.S.~{Booth}, eds., p.~380, Jan., 1993.

\bibitem{Kellerman1993}
K.I.~{Kellermann}, \emph{{The cosmological deceleration parameter estimated
  from the angular-size/redshift relation for compact radio sources}},
  \href{https://doi.org/10.1038/361134a0}{\emph{\nat} {\bfseries 361} (1993)
  134}.

\bibitem{Gurvits1994}
L.I.~{Gurvits}, \emph{{Apparent Milliarcsecond Sizes of Active Galactic Nuclei
  and the Geometry of the Universe}},
  \href{https://doi.org/10.1086/173999}{\emph{\apj} {\bfseries 425} (1994)
  442}.

\bibitem{Homan2000}
D.C.~{Homan} and J.F.C.~{Wardle}, \emph{{Direct Distance Measurements to
  Superluminal Radio Sources}},
  \href{https://doi.org/10.1086/308884}{\emph{\apj} {\bfseries 535} (2000) 575}
  [\href{https://arxiv.org/abs/astro-ph/0001298}{{\ttfamily
  astro-ph/0001298}}].

\bibitem{Cao2019_curvature}
S.~{Cao}, J.~{Qi}, M.~{Biesiada}, X.~{Zheng}, T.~{Xu}, Y.~{Pan} et~al.,
  \emph{{Milliarcsecond compact structure of radio quasars and the geometry of
  the Universe}},
  \href{https://doi.org/10.1016/j.dark.2019.100274}{\emph{Physics of the Dark
  Universe} {\bfseries 24} (2019) 100274}.

\bibitem{Cao2017_rulers}
S.~{Cao}, X.~{Zheng}, M.~{Biesiada}, J.~{Qi}, Y.~{Chen} and Z.-H.~{Zhu},
  \emph{{Ultra-compact structure in intermediate-luminosity radio quasars:
  building a sample of standard cosmological rulers and improving the dark
  energy constraints up to z 3}},
  \href{https://doi.org/10.1051/0004-6361/201730551}{\emph{\aap} {\bfseries
  606} (2017) A15} [\href{https://arxiv.org/abs/1708.08635}{{\ttfamily
  1708.08635}}].

\bibitem{Cohen1988}
M.H.~{Cohen}, P.D.~{Barthel}, T.J.~{Pearson} and J.A.~{Zensus},
  \emph{{Expanding Quasars and the Expansion of the Universe}},
  \href{https://doi.org/10.1086/166352}{\emph{\apj} {\bfseries 329} (1988) 1}.

\bibitem{Kellerman2004}
K.I.~{Kellermann}, M.L.~{Lister}, D.C.~{Homan}, R.C.~{Vermeulen}, M.H.~{Cohen},
  E.~{Ros} et~al., \emph{{Sub-Milliarcsecond Imaging of Quasars and Active
  Galactic Nuclei. III. Kinematics of Parsec-scale Radio Jets}},
  \href{https://doi.org/10.1086/421289}{\emph{\apj} {\bfseries 609} (2004) 539}
  [\href{https://arxiv.org/abs/astro-ph/0403320}{{\ttfamily
  astro-ph/0403320}}].

\bibitem{Jorstad2017}
S.G.~{Jorstad}, A.P.~{Marscher}, D.A.~{Morozova}, I.S.~{Troitsky}, I.~{Agudo},
  C.~{Casadio} et~al., \emph{{Kinematics of Parsec-scale Jets of Gamma-Ray
  Blazars at 43 GHz within the VLBA-BU-BLAZAR Program}},
  \href{https://doi.org/10.3847/1538-4357/aa8407}{\emph{\apj} {\bfseries 846}
  (2017) 98} [\href{https://arxiv.org/abs/1711.03983}{{\ttfamily 1711.03983}}].

\bibitem{Frey2015}
S.~{Frey}, Z.~{Paragi}, J.O.~{Fogasy} and L.I.~{Gurvits}, \emph{{The first
  estimate of radio jet proper motion at z > 5}},
  \href{https://doi.org/10.1093/mnras/stu2294}{\emph{\mnras} {\bfseries 446}
  (2015) 2921} [\href{https://arxiv.org/abs/1410.8101}{{\ttfamily 1410.8101}}].

\bibitem{Perger2018}
K.~{Perger}, S.~{Frey}, K.{\'E}.~{Gab{\'a}nyi}, T.~{An}, S.~{Britzen},
  H.-M.~{Cao} et~al., \emph{{Constraining the radio jet proper motion of the
  high-redshift quasar J2134-0419 at z = 4.3}},
  \href{https://doi.org/10.1093/mnras/sty837}{\emph{\mnras} {\bfseries 477}
  (2018) 1065} [\href{https://arxiv.org/abs/1803.10452}{{\ttfamily
  1803.10452}}].

\bibitem{An2020_mu_z}
T.~{An}, Y.~{Zhang} and S.~{Frey}, \emph{{A method for checking high-redshift
  identification of radio AGNs}},
  \href{https://doi.org/10.1093/mnras/staa2132}{\emph{\mnras} {\bfseries 497}
  (2020) 2260} [\href{https://arxiv.org/abs/2007.09038}{{\ttfamily
  2007.09038}}].

\bibitem{Wiik2001}
K.~{Wiik} and E.~{Valtaoja}, \emph{{The geometry of the universe from high
  resolution VLBI data of AGN shocks}},
  \href{https://doi.org/10.1051/0004-6361:20000285}{\emph{\aap} {\bfseries 366}
  (2001) 1061}.

\bibitem{Hodgson2020}
J.A.~{Hodgson}, B.~{L'Huillier}, I.~{Liodakis}, S.-S.~{Lee} and A.~{Shafieloo},
  \emph{{Using variability and VLBI to measure cosmological distances}},
  \href{https://doi.org/10.1093/mnrasl/slaa051}{\emph{\mnras} {\bfseries 495}
  (2020) L27} [\href{https://arxiv.org/abs/2003.10278}{{\ttfamily
  2003.10278}}].

\bibitem{Jorstad2016}
S.~{Jorstad} and A.~{Marscher}, \emph{{The VLBA-BU-BLAZAR Multi-Wavelength
  Monitoring Program}},
  \href{https://doi.org/10.3390/galaxies4040047}{\emph{Galaxies} {\bfseries 4}
  (2016) 47}.

\bibitem{Verde2019}
L.~{Verde}, T.~{Treu} and A.G.~{Riess}, \emph{{Tensions between the early and
  late Universe}},
  \href{https://doi.org/10.1038/s41550-019-0902-0}{\emph{Nature Astronomy}
  {\bfseries 3} (2019) 891} [\href{https://arxiv.org/abs/1907.10625}{{\ttfamily
  1907.10625}}].

\bibitem{Hodgson2023}
J.A.~{Hodgson}, B.~{L'Huillier}, I.~{Liodakis}, S.-S.~{Lee} and A.~{Shafieloo},
  \emph{{Estimating the feasibility of 'standard speed-gun' distances}},
  \href{https://doi.org/10.1093/mnrasl/slad007}{\emph{\mnras} {\bfseries 521}
  (2023) L44} [\href{https://arxiv.org/abs/2301.06252}{{\ttfamily
  2301.06252}}].

\bibitem{Readhead1994}
A.C.S.~{Readhead}, \emph{{Equipartition Brightness Temperature and the Inverse
  Compton Catastrophe}}, \href{https://doi.org/10.1086/174038}{\emph{\apj}
  {\bfseries 426} (1994) 51}.

\bibitem{Refsdal1964}
S.~{Refsdal}, \emph{{On the possibility of determining Hubble's parameter and
  the masses of galaxies from the gravitational lens effect}},
  \href{https://doi.org/10.1093/mnras/128.4.307}{\emph{\mnras} {\bfseries 128}
  (1964) 307}.

\bibitem{TreuMarshall2016}
T.~{Treu} and P.J.~{Marshall}, \emph{{Time delay cosmography}},
  \href{https://doi.org/10.1007/s00159-016-0096-8}{\emph{\aapr} {\bfseries 24}
  (2016) 11} [\href{https://arxiv.org/abs/1605.05333}{{\ttfamily 1605.05333}}].

\bibitem{Barnacka2015}
A.~{Barnacka}, M.J.~{Geller}, I.P.~{Dell'Antonio} and W.~{Benbow},
  \emph{{Resolving the High-energy Universe with Strong Gravitational Lensing:
  The Case of PKS 1830-211}},
  \href{https://doi.org/10.1088/0004-637X/809/1/100}{\emph{\apj} {\bfseries
  809} (2015) 100} [\href{https://arxiv.org/abs/1504.05210}{{\ttfamily
  1504.05210}}].

\bibitem{Spingola2016}
C.~{Spingola}, D.~{Dallacasa}, M.~{Orienti}, M.~{Giroletti}, J.P.~{McKean},
  C.C.~{Cheung} et~al., \emph{{Radio follow-up of the {\ensuremath{\gamma}}-ray
  flaring gravitational lens JVAS B0218+357}},
  \href{https://doi.org/10.1093/mnras/stw136}{\emph{\mnras} {\bfseries 457}
  (2016) 2263} [\href{https://arxiv.org/abs/1601.03591}{{\ttfamily
  1601.03591}}].

\bibitem{Biggs2018a}
A.D.~{Biggs} and I.W.A.~{Browne}, \emph{{A revised lens time delay for JVAS
  B0218+357 from a reanalysis of VLA monitoring data}},
  \href{https://doi.org/10.1093/mnras/sty565}{\emph{\mnras} {\bfseries 476}
  (2018) 5393} [\href{https://arxiv.org/abs/1802.10088}{{\ttfamily
  1802.10088}}].

\bibitem{Biggs2018b}
A.D.~{Biggs}, \emph{{The time delay of JVAS B1030+074 from VLA polarization
  monitoring}}, \href{https://doi.org/10.1093/mnras/sty2098}{\emph{\mnras}
  {\bfseries 481} (2018) 1000}
  [\href{https://arxiv.org/abs/1808.00757}{{\ttfamily 1808.00757}}].

\bibitem{Biggs2021}
A.D.~{Biggs}, \emph{{The time delay of CLASS B1600+434 from VLA multi-frequency
  and polarization monitoring}},
  \href{https://doi.org/10.1093/mnras/stab1444}{\emph{\mnras} {\bfseries 505}
  (2021) 2610} [\href{https://arxiv.org/abs/2105.07905}{{\ttfamily
  2105.07905}}].

\bibitem{Deller2011}
A.T.~{Deller}, W.F.~{Brisken}, C.J.~{Phillips}, J.~{Morgan}, W.~{Alef},
  R.~{Cappallo} et~al., \emph{{DiFX-2: A More Flexible, Efficient, Robust, and
  Powerful Software Correlator}},
  \href{https://doi.org/10.1086/658907}{\emph{\pasp} {\bfseries 123} (2011)
  275} [\href{https://arxiv.org/abs/1101.0885}{{\ttfamily 1101.0885}}].

\bibitem{Morgan2011}
J.S.~{Morgan}, F.~{Mantovani}, A.T.~{Deller}, W.~{Brisken}, W.~{Alef},
  E.~{Middelberg} et~al., \emph{{VLBI imaging throughout the primary beam using
  accurate UV shifting}},
  \href{https://doi.org/10.1051/0004-6361/201015775}{\emph{\aap} {\bfseries
  526} (2011) A140} [\href{https://arxiv.org/abs/1302.6040}{{\ttfamily
  1302.6040}}].

\bibitem{Radcliffe2016}
J.F.~{Radcliffe}, M.A.~{Garrett}, R.J.~{Beswick}, T.W.B.~{Muxlow},
  P.D.~{Barthel}, A.T.~{Deller} et~al., \emph{{Multi-source self-calibration:
  Unveiling the microJy population of compact radio sources}},
  \href{https://doi.org/10.1051/0004-6361/201527980}{\emph{\aap} {\bfseries
  587} (2016) A85} [\href{https://arxiv.org/abs/1601.04452}{{\ttfamily
  1601.04452}}].

\bibitem{Muxlow2020}
T.W.B.~{Muxlow}, A.P.~{Thomson}, J.F.~{Radcliffe}, N.H.~{Wrigley},
  R.J.~{Beswick}, I.~{Smail} et~al., \emph{{The e-MERGE Survey (e-MERLIN Galaxy
  Evolution Survey): overview and survey description}},
  \href{https://doi.org/10.1093/mnras/staa1279}{\emph{\mnras} {\bfseries 495}
  (2020) 1188} [\href{https://arxiv.org/abs/2005.02407}{{\ttfamily
  2005.02407}}].

\bibitem{Jackson2011}
N.~{Jackson}, \emph{{The Faintest Radio Source Yet: Expanded Very Large Array
  Observations of the Gravitational Lens SDSS J1004+4112}},
  \href{https://doi.org/10.1088/2041-8205/739/1/L28}{\emph{\apjl} {\bfseries
  739} (2011) L28} [\href{https://arxiv.org/abs/1105.2141}{{\ttfamily
  1105.2141}}].

\bibitem{Badole2020}
S.~{Badole}, N.~{Jackson}, P.~{Hartley}, D.~{Sluse}, H.~{Stacey} and
  H.~{Vives-Arias}, \emph{{VLA and ALMA observations of the lensed radio-quiet
  quasar SDSS J0924+0219: a molecular structure in a 3 {\ensuremath{\mu}}Jy
  radio source}}, \href{https://doi.org/10.1093/mnras/staa1488}{\emph{\mnras}
  {\bfseries 496} (2020) 138}
  [\href{https://arxiv.org/abs/2005.13612}{{\ttfamily 2005.13612}}].

\bibitem{Deller2014}
A.T.~{Deller} and E.~{Middelberg}, \emph{{mJIVE-20: A Survey for Compact mJy
  Radio Objects with the Very Long Baseline Array}},
  \href{https://doi.org/10.1088/0004-6256/147/1/14}{\emph{\aj} {\bfseries 147}
  (2014) 14} [\href{https://arxiv.org/abs/1310.8191}{{\ttfamily 1310.8191}}].

\bibitem{Shabala2017}
S.S.~{Shabala}, A.~{Deller}, S.~{Kaviraj}, E.~{Middelberg}, R.J.~{Turner},
  Y.S.~{Ting} et~al., \emph{{Delayed triggering of radio active galactic nuclei
  in gas-rich minor mergers in the local Universe}},
  \href{https://doi.org/10.1093/mnras/stw2536}{\emph{\mnras} {\bfseries 464}
  (2017) 4706} [\href{https://arxiv.org/abs/1608.04178}{{\ttfamily
  1608.04178}}].

\bibitem{Spingola2019_mjive}
C.~{Spingola}, J.P.~{McKean}, M.~{Lee}, A.~{Deller} and J.~{Moldon}, \emph{{A
  novel search for gravitationally lensed radio sources in wide-field VLBI
  imaging from the mJIVE-20 survey}},
  \href{https://doi.org/10.1093/mnras/sty3189}{\emph{\mnras} {\bfseries 483}
  (2019) 2125} [\href{https://arxiv.org/abs/1811.09152}{{\ttfamily
  1811.09152}}].

\bibitem{Burke-Spolaor2015}
S.~{Burke-Spolaor}, ``{Followup of Dual AGN Candidates from the mJIVE-20
  Survey}.'' Keck Observatory Archive LRIS, id.C271LA, Feb., 2015.

\bibitem{Rezaei2022}
S.~{Rezaei}, J.P.~{McKean}, M.~{Biehl} and A.~{Javadpour}, \emph{{DECORAS:
  detection and characterization of radio-astronomical sources using deep
  learning}}, \href{https://doi.org/10.1093/mnras/stab3519}{\emph{\mnras}
  {\bfseries 510} (2022) 5891}
  [\href{https://arxiv.org/abs/2109.09077}{{\ttfamily 2109.09077}}].

\bibitem{Janssen2019}
M.~{Janssen}, ``{rPICARD: Radboud PIpeline for the Calibration of high Angular
  Resolution Data}.'' Astrophysics Source Code Library, record ascl:1905.015,
  May, 2019.

\end{thebibliography}\endgroup

\documentclass[a4paper,11pt]{article}

\usepackage{pos}

\title{Cosmological studies with VLBI}

\author*[a]{Cristiana Spingola}

\affiliation[a]{INAF $-$ Istituto di Radioastronomia,\\
 Via Gobetti 101, I$-$40129, Bologna, Italy}

\emailAdd{spingola@ira.inaf.it}

\abstract{Current cosmological controversies can be solved if a sufficient level of precision is achieved by observations. Future surveys with the next generation of telescopes will offer significantly improved depth and angular resolution with respect to existing observations, opening the so-called “era of precision cosmology”.
But, that era can be considered already started at the radio wavelengths with Very Long Baseline Interferometry (VLBI). In this paper, we give an overview on how VLBI is contributing to some open questions in contemporary cosmology by reaching simultaneously the largest distances and the smallest scales.}

\FullConference{%
  15th European VLBI Network Mini-Symposium and Users' Meeting (EVN2022)\\
  11-15 July 2022\\
  University College Cork, Ireland
}

\begin{document} 
\maketitle


\section{Introduction}

In the current standard cosmological framework $\Lambda$CDM (Dark Energy Cold Dark Matter), simple gas (mostly H and He) evolves into highly complex galaxies, which are gravitationally bound systems made of metal enriched gas, stars, dust, supermassive black holes, that can become active and have a strong effect on the surrounding medium. All of these components are believed to be embedded in a dark matter halo \cite{Cimatti}. The cosmological model expects that the growth of these structures is driven by the dark matter component and happens in an expanding Universe, as known for almost a century \cite{Hubble1929} but still controversial \cite{Riess2020}. We are in an era where we can approach cosmology from an observational rather than a theoretical standpoint. The $\Lambda$CDM model has been successful at explaining the observations of the large-scale structure of the Universe, but it presents some issues on the small (galactic and sub-galactic) scales \cite{Bullock2017}. To clearly assess if such small-scale problems affect \textsl{all} galaxies, it is mandatory to robustly characterize also high redshift (high-z) galaxies. To do so, a primary requirement is high angular resolution and deep observations.

Current VLBI arrays are ideal to perform such tests, as they can provide milliarcsecond (mas) and sub-mas angular resolution (at high frequency) and $\mu$Jy beam$^{-1}$ sensitivity observations. However, if we want to use VLBI arrays, the selected targets must emit radiation at the radio wavelengths. Among the possible sources to investigate the small-scales of high-z galaxies there are active galactic nuclei (AGN). In particular, those AGN that shows collimated jets ("jetted AGN") can reach extremely large luminosities because of the effect of Doppler boosting \cite{Padovani2017}. Therefore, their radio emission can be detected up to the highest redshifts, making them ideal targets for studying the most distant Universe. Since jetted AGN can be cosmologically distant sources, some of them could be \textsl{strongly gravitationally lensed} (see \cite{Treu2010} for a review on strong lensing). The magnification and amplification due to the presence of a foreground (lensing) massive and concentrated object can make possible the detection also of the faintest jetted AGN in the form of multiple  and distorted images. Thus, strong lensing can provide an unbiased view on the jetted AGN population across the cosmic time.  Both lensed and "unlensed" jetted AGN give valuable observational constraints that can directly test  the standard model.


In this paper, we describe the contribution of VLBI to some current open issues in cosmology related to AGN and their surrounding medium (Sec.~\ref{Sec:2}), dark matter (Sec.~\ref{Sec:3}) and dark energy (Sec.~\ref{Sec:4}). Finally, we conclude by discussing the role of current and future wide-field VLBI surveys for observational cosmology (Sec.~\ref{Sec:5}).

\section{AGN and molecular gas at high redshift}\label{Sec:2}

How supermassive black holes (SMBHs) form and evolve is still unclear. The presence of SMBHs of masses of $10^{8-9}$ M$_{\odot}$ when the Universe was less than a Gyr old \cite{Banados2018, Belladitta2022} supports models of high mass BH seeds and super- (or even hyper-) Eddington accretion events. Some models assume an initial phase dominated by BH-BH mergers lasting up to $z\sim6$. After this epoch, the accretion of the resulting SMBHs is believed to be driven by inflowing of their surrounding medium (e. g., \cite{Piana2021}). From an observational point of view, jetted AGN have been found in gas-rich environments, which supports some of these models and at the same time suggests that mergers could be the triggering of radio-loudness in AGN \cite{Chiaberge2015, Gao2020}. 
Therefore, observations of the molecular gas surrounding jetted AGN can be a powerful method to test both SMBHs formation models and the activation of the radio-loudness across the cosmic time.

Very recently, the James Webb Space Telescope (JWST) has opened a new frontier for the study of high redshift galaxies, with a large number of candidates \textsl{during} the reionization epoch (at photometric redshifts $z>7$, e.g. \cite{Whitler2023} and references therein). The observations with JWST are suggesting even older and more massive SMBHs than expected, which require complex super-Eddington growth \cite{Pacucci2022, Volonteri2023}. 
VLBI observations of jetted AGN are currently reported up to $z=6.82$ \cite{Momjian2021}, finding jets extended from a few to several hundreds of parsecs \cite{Momjian2008, Frey2010, Frey2011, Frey2013, Cao2014, Gabanyi2015, Zhang2017,Cao2017, Wang2017, Momjian2018, Frey2018, Perger2019, Spingola2020, An2020, Zhang2021, Gabanyi2021, Perger2021, Zhang2022, Krezinger2022}. 
As for the molecular gas, at high-z detections of serveral J transitions of the CO and C$^+$ emission lines have been reported in the radio/mm bands \cite{Walter2014, Aravena2016a, Aravena2016b,  Decarli2016, Riechers2019, Spingola2020_gas, Bischetti2021}, which revealed the presence of massive molecular gas reservoirs around SMBHs, in agreement with the predictions from hydrodynamical simulations \cite[e.g.,][]{DiMatteo2008,Dave2020}. 
Moreover, detailed analysis of the profile of these emission line showed evidence for on-going mergers in gas-rich radio-loud AGN at $z>6$ \cite{Khusanova2022}. However, these results have been obtained for the brightest high-z sources and these sources were only partially spatially resolved. In this context, the magnification due to strong lensing is really powerful for detecting the low surface brightness emission of the low-J transitions of CO at high redshift. This boost in spatial resolution and sensitivity revealed both mergers (complex velocity fields) but also molecular gas disks (ordered rotation) in jetted AGN at high redshift \cite{Riechers2008a, Riechers2008b, Riechers2011,Deane2013, Paraficz2018, Spingola2020_gas}.

As an intermediate stage of the merger process, we expect to observe binary (or multiple) SMBH systems \cite{Begelman1980, DiMatteo2005, Volonteri2021}.  The interesting multiple SMBHs systems for the future pulsar timing arrays and the Laser Interferometer Space Antenna are those at \textsl{small separation}, as they will be the emitters of gravitational waves in the $\mu$Hz--nHz regime \citep{BurkeSpolaor2019, Colpi2019}. The fraction of the active pairs of SMBHs (binary/dual AGN) is very low in the local Universe, but it increases up to a few percent at high redshift \citep{DeRosa2019}. Hence, ideally, the search for close pairs of SMBHs should be performed at $z>1$. Only VLBI can spatially resolve such systems on parsec scales even at large distances. Indeed, using VLBI observations several pairs of AGN have been found so far \citep{Rodriguez2006, Bondi2010, Burke-Spolaor2011, An2013, Deane2014, Gabanyi2016}.

The combination of strong lensing and VLBI has been proven to be a powerful method to unveil dual AGN candidates at high redshift and small separation. Three cases have been reported so far:  MG~B2016+112 at $z=3.273$ \citep{Spingola2019_mg2016, Schwartz2021}, PS J1721+8842 at $z=2.37$ \citep{Mangat2021} and PKS~1830-211 at $z=2.5$ \citep{Nair2005}. As first suggested by \citep{Spingola2019_mg2016}, the detection of even a single strongly lensed binary AGN implies a larger population of AGN pairs in the early Universe than first thought. However, such estimate is highly limited by the very small number statistics, and these lensed (and unlensed) candidates should also be confirmed. 
Nevertheless, VLBI provides currently the only way to find the closest AGN pairs, hence probing the bound systems and the final merging stages at the largest distances (as summarized in Fig. 7 of \citep{Spingola2022}).

\section{Dark matter distribution at high-z}\label{Sec:3}


Among the current small-scale issues of the $\Lambda$CDM model, the most persistent ones are the so-called \textsl{missing satellite} and the \textsl{cusp-core} problems \cite{Bullock2017, Delpopolo2021}. Both these observational-theoretical conflicts could be solved by questioning the cold dark matter paradigm and assuming an alternative ("warm") dark matter particle \citep{Lovell2012}. 
Gravitational lensing provides a direct way to both find the expected satellites orbiting around massive galaxies and infer the inner mass density profile to the the cusp-core problem. In particular, the strong lensing systems showing many extended images in radial and tangential directions are ideal for such investigations. One can detect the sub-halos associated with a lensing galaxy or along the line of sight by means of "anomalies" in the flux, position or time delay in the lensed images. 
Such "anomalies" arise because the presence of a sub-halo that locally changes the magnification of that image, hence its properties (in particular flux and position) will be different from what expected for a smooth mass density distribution (i.e. without sub-halos). For instance, \citep{Hartley2019} recently detected mas-level positional anomalies in the lensed images positions inferred by VLBI observations in HS 0810+2554. This system shows eight extended lensed images of a background AGN with intrinsic integrated flux density of only 880 nJy.
Given the angular scale of these astrometric perturbations, they could be due to a population of low-mass (10$^{5-7}$ M$_{\odot}$) dark matter dominated sub-halos. 
Milliarcsecond-level astrometric anomalies were found also using global-VLBI observations of the lensing system MG J0751+2716, which shows compact and extended lensed images forming giant gravitational arcs \citep{Spingola2018}. Therefore, also for this system it is possible that the observed astrometric anomalies are due to low-mass sub-halos (see also \citep{Biggs2004, McKean2007, More2009, Hsueh2016, Hsueh2017, Stacey2020}). The inner density profile of MG J0751+2716 was inferred at high precision, thanks to the many constraints from the VLBI-detected lensed images. The inner mass density profile has been found to be steeper than isothermal: such steepening could be caused by the dissipative process due to \textsl{in situ} star formation, which leads to an an increased stellar content in the central regions of the lensing galaxy. Therefore, it is possible that the complex processes involving \textsl{baryonic} matter might be responsible for the current tensions regarding the small scales of galaxies. Only VLBI observations of extended gravitational arcs can test such complex models, as demonstrated by \citep{Powell2021, Powell2022} for testing the origin of the astrometric anomalies in MG J0751+2716.

Despite the advantages of using gravitational lensing and VLBI to study astrometric anomalies at mas and sub-mas level, we are currently limited by the small number of VLBI-detected systems. More strongly lensed jetted AGN observed with VLBI are needed to provide a conclusive interpretation on these dark matter studies.

\section{Cosmological parameters from jetted AGN observed with VLBI}
\label{Sec:4}


The research field of using VLBI observations of AGN to infer cosmological parameters (such as $q_0$ and $H_0$) has been pioneered by \citep{Gurvits1993, Kellerman1993}. They proposed a characteristic angular size $\theta$ for compact jetted AGN to build a cosmic distance ladder. For a cosmological rod of intrinsic length $l_m$, the $\theta$-z relation can be written as $\theta(z) = l_m/ D_A(z)$, where $D_A$ is the angular diameter distance, which depends on the cosmological parameters. This method was applied for the first time using $\sim340$ VLBI observations of AGN with known redshift by \cite{Gurvits1994}. With these observations they built a $\theta-z$ plane up to $z\simeq4$ and could test different cosmological models (namely, different values of $q_0$), despite the large uncertainty given by the limited number of sources.  Also, early work of \cite{Homan2000} proposed to measure the distance to jetted AGN by combining the observed proper motion, Doppler factor derived from equipartition between the energy density in magnetic and particles fields, and aspect ratio of the (superluminal) jet component. Recently \cite{Cao2019_curvature} proposed to use VLBI-detected AGN also to measure the cosmic curvature ($\Omega_k$), finding no significant deviations from a flat Universe.
The caveat about these methods is whether it is possible to use jetted AGN as standard rods. For this reason, a lot of research has been done to find the best jetted AGN "category" or some of their components for this cosmological application (see \citep{Cao2017_rulers} and references therein).

Another methodology to test cosmological models is to measure jet proper motions ($\mu$) across the cosmic time, as $\mu$ depends on the cosmological parameters \citep{Cohen1988, Kellerman2004, Jorstad2017}. The difficulty of this method is to obtain measurements of proper motions at high-z, as such measurements require long time baselines and high signal-to-noise ratios. Nevertheless, such measurements have been obtained up to $z\simeq 5.5$ (e.g., \citep{Frey2015, Perger2018, An2020, Zhang2022} and references therein). Recently, a reverse application of the $\mu$-z plot has been proposed, namely for deriving the redshift of a radio-loud AGN by adopting the standard cosmological model \citep{An2020_mu_z}. This method will be particularly important for wide-field surveys, where there are AGN candidates lacking spectroscopic redshift.

A method to estimate $H_0$ is based on the variability of jetted AGN \cite{Wiik2001} and has been recently used by
\cite{Hodgson2020} on the Boston University blazar monitoring  programme \cite{Jorstad2016}. The approach is based on light travel-time ("causality") arguments, and can be extended from low- to high-z AGN. The angular diameter distance $D_A$ can be also expressed as $D_A(z) = c \Delta t / \theta_{\rm VLBI} (1+z)$, where $\Delta t$ is the variability  timescale and $\theta_{\rm VLBI}$ is the size of size of the variable region. By applying this method on the low-z AGN 3C~84, they infer a value for $H_0$ with a similar precision of other "late Universe methods" applied to samples of sources (summarized in \citep{Verde2019}), demonstrating  that this approach is a competitive cosmological probe. Recently, \cite{Hodgson2023} revised the most important issue of this methodology, which is the determination of the Doppler factor, which becomes more difficult to obtain for high-z jetted AGN. The authors propose a method to derive the Doppler factor that is independent from the typical equipartition and inverse Compton brightness temperature limits \cite{Readhead1994}.

Finally, gravitational time delays observed in strong lensing systems provide a one-step measurement for $H_0$ \citep{Refsdal1964}, because the time-delay distance is inversely proportional to $H_0$ \cite{TreuMarshall2016}. Even if time delays between the multiple images can be estimated also using a spatially unresolved light curve (e.g., \cite{Barnacka2015}), at the radio wavelengths i) VLBI can easily resolve the lensed images (typically separated by arcseconds), ii) AGN show intrinsic time variability on several time scales and iii) there is not the additional challenge of microlensing, as the (background) jetted AGN are typically extended several parsecs in size. Despite the powerful combination of VLBI and gravitational time delays, there has not been any time delay measured using VLBI yet. This is mostly due to the high cadence of the observing epochs required and the relatively short periods for such monitoring programs available  (e.g., \citep{Spingola2016}).  Instead, monitoring programs with the Very Large Array have been proven to be quite effective for estimating gravitational time delays also by innovatively using  polarization \citep{Biggs2018a, Biggs2018b, Biggs2021}.

\section{Wide-field VLBI for cosmological probes and conclusions}\label{Sec:5}

Wide-field VLBI is now possible routinely. In the last $\sim10$ years, impressive progress has been done to make this possible, and in particular the development of multi-phase centre correlation (i.e., the entire primary beam of VLBI experiment can be imaged \cite{Deller2011, Morgan2011}) and multi-source self-calibration \cite{Radcliffe2016}. The progress in wide-field VLBI is timely, as there is the need to complement current (and future) surveys with radio observations at sub-arcsec angular resolution. At the same time, the exquisite sensitivity that wide-field VLBI surveys can reach now (e.g., $1.5\,\mu$Jy beam$^{-1}$ in the \textsl{e}MERGE survey \cite{Muxlow2020}) enables us to study at high image fidelity the faintest AGN population, which only the lensing magnification revealed so far \cite{Jackson2011, Hartley2019, Badole2020}.
Hence, wide-field VLBI surveys can detect the different radio sources that are most suitable for each of the cosmological applications described in this paper. For instance, the mJIVE-20 survey at 1.4 GHz \citep{Deller2014} has been used to search for radio-loud AGN in gas-rich mergers \cite{Shabala2017}, to find strongly lensed jetted AGN at small angular separation \citep{Spingola2019_mjive} and for the hunt for AGN pairs \citep{Burke-Spolaor2015}. The same wide-field VLBI survey has been also used for developing tools to identify and characterize VLBI sources with deep learning \cite{Rezaei2022}. Such tools together with pipelines for calibrating large VLBI datasets (e.g., \cite{Janssen2019}) are crucial for current and future wide-field VLBI surveys.


In summary, VLBI can provide precise and independent cosmological probes. In this paper, we focused on how VLBI observations of AGN (lensed and unlensed) can test the current cosmological model. Future surveys with the Square Kilometer Array (SKA) and the follow-ups with SKA-VLBI will offer significantly improved sensitivity. In the future,  we will not be limited by the small number statistics anymore, but we will also need synergy with multi-wavelength instruments to more robustly provide competitive precision on the estimate of the cosmological parameters.


\bibliographystyle{JHEP}
\bibliography{refs}

\end{document}